OpenStudy (anonymous):

How to solve. 6/(k^2-6k)=1/(k-6)+(k-5)/k

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

jhonyy9 (jhonyy9):

6/(k(k-6))=1/(k-6)+(k-5)/k --- 6=k+(k-6)(k-5) --- 6=k+k*2-11k+30 --- k*2-10k+24=0

jhonyy9 (jhonyy9):

solutions will be k=6 and k=4